OpenStudy (anonymous):

Solve: |2x - 7| < 13

OpenStudy (anonymous):

Case I 2x - 7 < 13 2x < 20 x < 10 case II - (2x - 7) < 13 2x - 7 > -13 2x > -6 x > -3

OpenStudy (anonymous):

solve it as: 2x - 7=13 Note: The inequality only changes its sign, if a number with "- minus" sign is shifted to other side of "= equal" sign

OpenStudy (anonymous):

2x + 7 =< 13 2x = 13 - 7..........By transporting + 7 to the other side makes it - 7. 2x = 6.................By subtracting 13 from 7 the result is 6 as 7 + 6 = 13. x = 6/2................By transporting 2 X x to the other side makes it 6 / 2. x = >-3...................By dividing 6 and 2 result is 3 as 2 X 3 = 6.